# Tele-LLMs: A Series of Specialized Large Language Models for Telecommunications

### v2

@@ Abstract.
The emergence of large language models (LLMs) has significantly impacted various fields, from natural language processing to sectors like medicine and finance. However, despite their rapid proliferation, the applications of LLMs in telecommunications remain limited, often relying on general-purpose models that lack domain-specific specialization. This lack of specialization results in underperformance, particularly when dealing with telecommunications-specific technical terminology and their associated mathematical representations. This paper addresses this gap by first creating and disseminating Tele-Data, a comprehensive dataset of telecommunications material curated from relevant sources, and Tele-Eval, a large-scale question-and-answer dataset tailored to the domain. Through extensive experiments, we explore the most effective training techniques for adapting LLMs to the telecommunications domain, ranging from examining the division of expertise across various telecommunications aspects to employing parameter-efficient techniques. We also investigate how models of different sizes behave during adaptation and analyze the impact of their training data on this behavior. Leveraging these findings, we develop and open-source Tele-LLMs [^1] The Hugging Face links to both the datasets and Tele-LLMs can be found at https://github.com/Ali-maatouk/Tele-LLMs   This work was conducted during Kenny’s time at Yale University., the first series of language models ranging from 1B to 8B parameters, specifically tailored for telecommunications. Our evaluations demonstrate that these models outperform their general-purpose counterparts on Tele-Eval while retaining their previously acquired capabilities, thus avoiding the catastrophic forgetting phenomenon.
1. Introduction
Figure 1. Overall pipeline of the LLM adaptation to telecommunications.
@@ 1. Introduction
Large language models (LLMs) have recently marked a major breakthrough in natural language processing. Since 2022, these deep learning systems have proliferated rapidly, with numerous models released by tech giants, research institutions, and open-source communities (OpenAI, 2024; et al., 2023, 2024). Trained on vast text corpora, LLMs have demonstrated unprecedented capabilities in understanding context, generating human-like text, and performing reasoning tasks across diverse domains (Radford et al., 2019). These abilities have sparked interest from researchers and industry professionals to adopt and explore their potential applications across a wide variety of fields.
@@
Researchers in the telecommunications and networking domain are no exception, aiming to leverage LLMs’ potential across various tasks within the field. These applications include chatbots for engineers (Kotaru, 2023), network document analysis (Bariah et al., 2023b), network modeling and development (Maatouk et al., 2024), and wireless systems design (Du et al., 2023; Xu et al., 2024). The list of tasks being explored for potential LLM support continues to grow (Bariah et al., 2023a), with new research emerging rapidly to assess the limits and utility of these models in the telecommunications context.
@@
Thus far, applications of LLMs in the telecommunications domain have primarily involved prompting (Du et al., 2023; Maatouk et al., 2024), in-context learning (Kotaru, 2023), and task-specific fine-tuning (Zhou et al., 2024; Bariah et al., 2023b), utilizing either proprietary LLMs like OpenAI’s GPT (OpenAI, 2024) or local open-source generic LLMs like Meta’s LLaMA series (et al., 2024). However, proprietary LLMs raise privacy concerns, as they require sharing prompts and related data with the LLM owners. Additionally, they offer limited maneuverability and adaptability, as users do not have access to the model weights. Furthermore, despite their adaptability, general-purpose open-source LLMs lack specialization in telecommunications, as they encompass knowledge from diverse domains such as medicine, history, and law. This lack of specialization hinders their performance; in fact, it is well established that LLM applications in specific domains, from prompting to task-specific fine-tuning, perform better when using domain-specific LLMs rather than general ones (Gururangan et al., 2020).
@@
Given the above, in a multitude of domains, efforts have been made to create specialized open-source LLMs to push the performance limits of LLMs applications in those fields, such as medicine (Labrak et al., 2024), law (Colombo et al., 2024), and finance (Xie et al., 2023). However, to date, there are no domain-specific open-source models in the telecommunications literature. This stems from the particularities of the telecommunications domain, which includes heterogeneous material such as standards and scholarly work, a heavy reliance on complex equations, varying formatting in datasets, and the generally closed-source nature of work done by companies thus far with LLMs (Kotaru, 2023; Holm, 2021). Given the anticipated role of LLMs in telecommunications, this represents a significant gap in the field.
@@
Proposed Work. This work aims to address the aforementioned gap by introducing the first series of LLMs specialized for the telecommunications field. Our contribution extends beyond open-sourcing the models; we open-source every step of the framework required for this specialization. We also provide, through extensive experiments, key insights into the process of adapting LLMs to the telecommunications domain. This includes the techniques to be used and the different behaviors exhibited by various LLMs during adaptation, based on their sizes and overall tendencies.
@@
Second, leveraging Tele-Data, we use an LLM-based framework to create an evaluation dataset consisting of 750k question-and-answer (QnA) pairs, referred to as Tele-Eval. This dataset represents the first large-scale open-ended QnA dataset in the telecommunications field. It also includes references to the specific material from which each question was generated, allowing users to select questions from the material they are interested in and facilitating retrieval-augmented generation frameworks (Lewis et al., 2021). To ensure the relevance and quality of the questions generated by the LLM, we apply strict regular expressions and LLM-based filtering to exclude questions of local interest, such as simulation results or locally-defined content.
Third, we examine the viability of using parameter-efficient fine-tuning (PEFT) techniques to inject telecommunications knowledge into these models, demonstrating how this injection instead requires the full fine-tuning (FFT) paradigm. Based on this, we conduct experiments to determine the number of training epochs needed to maximize model performance and identify the point at which overfitting occurs during this adaptation. Throughout this exploration, we also examine how models of different sizes behave during the adaptation process and investigate how the training dynamics of this adaptation vary based on the original model’s training data. Furthermore, we assess the effectiveness of dividing this adaptation by training multiple specialized models focused on different aspects of telecommunications rather than one overall combined model. Our findings demonstrate that the latter approach is superior due to the transfer learning that occurs across these aspects.
Finally, our experiments culminate in the adaptation to telecommunications and open-sourcing of a series of LLMs, ranging from 1B to 8B parameters. Our series is based on Tinyllama-1.1B, Phi-1.5, Gemma-2B, and LLama-3-8B, and includes both the base models and their instruct-fine-tuned versions for chatbot applications. We evaluate these models by comparing them with their original versions through both quantitative analyses, using a set of proposed metrics, and qualitative assessments. The results highlight their superiority in the telecom domain, achieving an average of 25% relative improvement on Tele-Eval. Additionally, our findings demonstrate how smaller adapted models can rival larger general-purpose models on Tele-Eval. The evaluation also indicates that these models retain their original capabilities, effectively avoiding the catastrophic forgetting phenomenon. The overall adaptation pipeline is provided in Fig. 1.
2. Domain Adaptation
2.1. Background
@@
As demonstrated in (Çağatay Yıldız et al., 2024), if one intends to use an LLM solely for a specific domain, specializing the LLM in that domain is advantageous for any subsequent in-context learning or task-specific fine-tuning. This is because the probability distribution of tokens in the domain of interest can differ significantly from those in other domains internalized by the LLM. By adapting to this distribution, the LLM’s knowledge becomes more targeted to the domain, leading to transfer learning for subsequent usage stages.
The primary approach for domain adaptation is continual pretraining. This method involves further training the LLM on domain-specific corpora, allowing it to adapt its parameters to the field of interest (Gururangan et al., 2020). Continual pretraining has become the standard for LLM domain adaptation, as evidenced by its application in various fields such as medicine (Labrak et al., 2024), law (Colombo et al., 2024), and finance (Xie et al., 2023). In the following, we will explore the details of continual pretraining, with a particular emphasis on its application in the telecommunications domain.
Remark 1.
@@
2.2. Continual Pretraining
Given the scale of the training datasets of current LLMs (e.g., 15T tokens for LLama-3 (et al., 2024)), it is fair to assume that the language models available in the literature have been trained on all available public-source data. Therefore, these models have likely already encountered most of the publicly available telecommunications-related data. Nevertheless, the goal of continual pretraining is to adapt the LLM’s knowledge to this specific domain by re-exposing the model to these data. Specifically, consider an LLM trained on a corpus of tokens $\mathcal{D}$ drawn from a distribution $\mu$. This LLM is trained over the entire corpus $\mathcal{D}$ to minimize the cross-entropy loss function for tokens $\bm{x}=(x_{1},\ldots,x_{T})$ drawn from this corpus:
(1) $\min_{\bm{W}\in\mathbb{R}^{n}}\mathbb{E}_{\bm{x}\sim\mu}\left[\mathcal{L}_{\text{CE}}(\bm{x};\bm{W})\right]=-\mathbb{E}_{\bm{x}\sim\mu}\left[\sum_{t=1}^{T}\log P(x_{t}\mid\bm{x}_{1:t-1};\bm{W})\right]$
@@ 2.2. Continual Pretraining
(2) $\min_{\bm{W}\in\mathbb{R}^{n}}-\mathbb{E}_{\bm{x}\sim\sigma}\left[\sum_{t=1}^{T}\log P(x_{t}\mid\bm{x}_{1:t-1};\bm{W},\bm{W}_{0})\right].$
@@
2.3. Catastrophic Forgetting
Ideally, one would hope that such domain adaptation comes at no penalty. However, generally, a penalty is incurred. By shifting the focus from the general token distribution $\mu$ of the training dataset $\mathcal{D}$ to a new distribution $\sigma$, the model risks forgetting previously acquired knowledge. This risk depends on the size of the model, the size of the domain-specific dataset, and how different $\sigma$ is from $\mu$ (Luo et al., 2024). Addressing this issue is crucial; otherwise, one may end up with a telecom-knowledge proficient LLM but lose reasoning capabilities, coding abilities, and general English language understanding in the process. Later in this paper, specifically in Section 7, we will detail how we addressed this issue in our continual pretraining framework. With this in mind, the next step in our framework involves curating a comprehensive telecommunications corpus, which we will refer to as Tele-Data.
3. Tele-Data
@@
Table 1. Tele-Data division across categories.
@@
Table 2. Datasets filtering statistics.
Figure 2. Raw cross-entropy loss.
Figure 3. Cleaned cross-entropy loss.
Our collection of telecommunications material comprises four main sources: (1) scientific papers from arXiv, (2) 3GPP standards, (3) Wikipedia articles related to telecommunications, and (4) telecommunications-related websites extracted from Common Crawl dumps. This diverse range of sources ensures broad coverage of telecommunications knowledge and is essential for enabling the transfer of expertise from one aspect of telecommunications covered in one source to another, as will be shown later in Section 8.1. The detailed composition of our dataset is reported in Table 1, with LLama-3 tokenizer used to provide token counts as an example.
3.1. Arxiv
@@
To assess the quality of our filtering process, we randomly sampled 500 arXiv papers and manually annotated them to determine their relevance to the telecommunications and networking domain. To ensure a balance between relevant and irrelevant examples, we drew an equal number of samples from both the computer science and electrical engineering categories given that the latter is more likely to contain telecommunications material, while the former is more likely to include unrelated papers. We then evaluated both the precision and recall of the filtering process. As shown in Table 2, the process demonstrates high recall and moderate precision. High recall is particularly important, as it indicates that our filtering process captures virtually all relevant telecommunications material. The moderate precision can be attributed to the inclusion of papers from related domains. For instance, some papers investigating cloud computing topics were included. Since these additional domains are not orthogonal to telecommunications, they can provide transferable knowledge that may enhance the language model’s adaptation to the telecommunications domain. All in all, this demonstrates the effectiveness of our filtering process.
Cleaning. Following curation, we implement a comprehensive cleaning process for the papers, which includes: (1) removing comments, (2) flattening LaTeX sources, (3) substituting user-defined macros with standard LaTeX commands, (4) eliminating LaTeX native commands, and (5) standardizing citation formats and removing markup changes. We also remove figures and tables to focus on inline text and equations. Details on this procedure are provided in Appendix A. To assess the effectiveness of our cleaning process, we randomly selected 500 arXiv papers and derived a cross-entropy-based approach to evaluate cleanliness as detailed below. Let $\mathcal{X}=(\bm{x}^{1},\bm{x}^{2},\ldots,\bm{x}^{N})$ and $\mathcal{Y}=(\bm{y}^{1},\bm{y}^{2},\ldots,\bm{y}^{M})$ denote the non-overlapping chunks of $T$ tokens from the raw and cleaned datasets, respectively, where $N$ and $M$ are the total number of chunks. Next, let us define the cross-entropy loss for any token index $t\in\{1,\ldots,T\}$ in any chunk of either the raw or cleaned data using GPT-2 (Radford et al., 2019) as follows
@@
(3) $\mathcal{L}^{\text{clean}}_{\text{CE}}(y_{t}^{i}|\bm{y}_{1:t-1}^{i})=-\log P(y_{t}^{i}|y_{1:t-1}^{i};\bm{W}_{\text{GPT-2}}),\quad i=1,\ldots,M.$
GPT-2 was chosen because it was primarily trained on web content, excluding arXiv scientific papers. Therefore, the cross-entropy here serves as a proxy for how closely the text resembles generic online content rather than the more complex arXiv papers, thus providing an indication of the dataset’s cleanliness. With this in mind, we set $T$ to 1024, in line with GPT-2’s training data (Radford et al., 2019), and define the total set of cross-entropy losses for the raw and cleaned datasets as
@@
(4) $\mathcal{C}=\bigcup_{i=1}^{M}\{\mathcal{L}_{\text{CE}}(y_{t}^{i}|\bm{y}_{1:t-1}^{i}):t\in[1,1024]\},$

### v3

@@ Abstract
The emergence of large language models (LLMs) has significantly impacted various fields, from natural language processing to sectors like medicine and finance. However, despite their rapid proliferation, the applications of LLMs in telecommunications remain limited, often relying on general-purpose models that lack domain-specific specialization. This lack of specialization results in underperformance, particularly when dealing with telecommunications-specific technical terminology and their associated mathematical representations. This paper addresses this gap by first creating and disseminating Tele-Data, a comprehensive dataset of telecommunications material curated from relevant sources, and Tele-Eval, a large-scale question-and-answer dataset tailored to the domain. Through extensive experiments, we explore the most effective training techniques for adapting LLMs to the telecommunications domain, ranging from examining the division of expertise across various telecommunications aspects to employing parameter-efficient techniques. We also investigate how models of different sizes behave during adaptation and analyze the impact of their training data on this behavior. Leveraging these findings, we develop and open-source Tele-LLMs [^1] The Hugging Face links to both the datasets and Tele-LLMs can be found at https://github.com/Ali-maatouk/Tele-LLMs   This work was conducted during Kenny’s time at Yale University., the first series of language models ranging from 1B to 8B parameters, specifically tailored for telecommunications. Our evaluations demonstrate that these models outperform their general-purpose counterparts on Tele-Eval and telecommunications-related literature tasks while retaining their previously acquired capabilities, thus avoiding the catastrophic forgetting phenomenon.
I Introduction
Figure 1: Overall pipeline of the LLM adaptation to telecommunications.
@@ I Introduction
Large language models (LLMs) have recently marked a major breakthrough in natural language processing. Since 2022, these deep learning systems have proliferated rapidly, with numerous models released by tech giants, research institutions, and open-source communities [1, 2, 3]. Trained on vast text corpora, LLMs have demonstrated unprecedented capabilities in understanding context, generating human-like text, and performing reasoning tasks across diverse domains [4]. These abilities have sparked interest from researchers and industry professionals to adopt and explore their potential applications across a wide variety of fields.
@@
Researchers in the telecommunications and networking domain are no exception, aiming to leverage LLMs’ potential across various tasks within the field. These applications include chatbots for engineers [5], network document analysis [6], network modeling and development [7], and wireless systems design [8, 9]. The list of tasks being explored for potential LLM support continues to grow [10], with new research emerging rapidly to assess the limits and utility of these models in the telecommunications context.
@@
Thus far, applications of LLMs in the telecommunications domain have primarily involved prompting [8, 7], in-context learning [5], and task-specific fine-tuning [11, 6], utilizing either proprietary LLMs like OpenAI’s GPT [1] or local open-source generic LLMs like Meta’s LLaMA series [3]. However, proprietary LLMs raise concerns as they may require sharing prompts and related data with the LLM owners. Additionally, they offer limited maneuverability and adaptability since users do not have access to the model weights. Furthermore, despite their adaptability, general-purpose open-source LLMs lack specialization in telecommunications, as they encompass knowledge from diverse domains such as medicine, history, and law. This lack of specialization hinders their performance; in fact, it is well established that LLM applications in specific domains, from prompting to task-specific fine-tuning, perform better when using domain-specific LLMs rather than general ones [12].
@@
Given the above, in a multitude of domains, efforts have been made to create specialized open-source LLMs to push the performance limits of LLMs applications in those fields, such as medicine [13], law [14], and finance [15]. However, to date, there are no domain-specific open-source models in the telecommunications literature. This stems from the particularities of the telecommunications domain, which includes heterogeneous material such as standards and scholarly work, a heavy reliance on complex equations, varying formatting in datasets, and the generally closed-source nature of work done by companies and institutions thus far with LLMs [5, 16, 17]. Given the anticipated role of LLMs in telecommunications, this represents a significant gap in the field.
@@
Proposed Work. This work aims to address the aforementioned gap by introducing the first open-source series of LLMs specialized for the telecommunications field. Our contribution extends beyond open-sourcing the models; we open-source every step of the framework required for this specialization. We also provide, through extensive experiments, key insights into the process of adapting LLMs to the telecommunications domain. This includes the techniques to be used and the different behaviors exhibited by various LLMs during adaptation, based on their sizes and overall tendencies.
@@
Second, leveraging Tele-Data, we use an LLM-based framework to create an evaluation dataset consisting of 750k question-and-answer (QnA) pairs, referred to as Tele-Eval. This dataset represents the first large-scale open-ended QnA dataset in the telecommunications field. It also includes references to the specific material from which each question was generated, allowing users to select questions from the material they are interested in and facilitating retrieval-augmented generation frameworks [18]. To ensure the relevance and quality of the questions generated by the LLM, we apply strict regular expressions and LLM-based filtering to exclude questions of local interest, such as simulation results or locally-defined content.
Third, we examine the viability of using parameter-efficient fine-tuning (PEFT) techniques to inject telecommunications knowledge into these LLMs. Our findings show that, instead, full fine-tuning (FFT) is required. Based on this, we conduct experiments to determine the number of training epochs needed to maximize model performance and to identify when overfitting occurs. We also explore how models of different sizes behave during adaptation and how training dynamics vary depending on the original model’s data. Furthermore, we assess whether splitting the adaptation into multiple specialized models, each focused on different aspects of telecommunications, is more effective than a single, combined model. Our results demonstrate that the latter approach is superior due to the transfer learning that occurs across these aspects.
Finally, building on these findings, our work culminates in adapting and open-sourcing a series of telecommunications-specialized LLMs ranging from 1B to 8B parameters. Our series is based on TinyLlama-1.1B, Phi-1.5B, Gemma-2B, Gemma-2-2B, LLaMA-3.2-1B, LLaMA-3.2-3B, and LLaMA-3-8B, and includes both base models and their instruction-tuned versions for chatbot applications. To evaluate our telecom-adapted models, we first compare them against their original counterparts using Tele-Eval, employing both quantitative analysis—based on a set of proposed metrics—and qualitative assessment. Our results show an average relative improvement of 25% on Tele-Eval, thus demonstrating specialization for telecom knowledge while preserving the models’ original capabilities and avoiding catastrophic forgetting. Next, to further evaluate our telecom-adapted models, we curate downstream tasks related to telecom literature, such as citation prediction and recommendation. Our results show that the telecom-specialized models consistently outperform their original counterparts on these tasks. Moreover, when fine-tuned appropriately, they can rival larger general-purpose models like GPT-4o on these telecom literature applications. The overall adaptation pipeline is provided in Fig. 1.
II Domain Adaptation
II-A Background
@@
As demonstrated in [19], if one intends to use an LLM solely for a specific domain, specializing the LLM in that domain is advantageous for any subsequent in-context learning or task-specific fine-tuning. This is because the probability distribution of tokens in the domain of interest can differ significantly from those in other domains internalized by the LLM. By adapting to this distribution, the LLM’s knowledge becomes more targeted to the domain, leading to transfer learning for subsequent usage stages.
The primary approach for domain adaptation is continual pretraining. This method involves further training the LLM on domain-specific corpora, allowing it to adapt its parameters to the field of interest [12]. Continual pretraining has become the standard for LLM domain adaptation, as evidenced by its application in various fields such as medicine [13], law [14], and finance [15]. In the following, we will explore the details of continual pretraining, with a particular emphasis on its application in the telecommunications domain.
Remark 1
@@
II-B Continual Pretraining
Given the scale of the training datasets of current LLMs (e.g., 15T tokens for LLama-3 [3]), it is fair to assume that the language models available in the literature have been trained on all available public-source data. Therefore, these models have likely already encountered most of the publicly available telecommunications-related data. Nevertheless, the goal of continual pretraining is to adapt the LLM’s knowledge to this specific domain by re-exposing the model to these data. Specifically, consider an LLM trained on a corpus of tokens $\mathcal{D}$ drawn from a distribution $\mu$. This LLM is trained over the entire corpus $\mathcal{D}$ to minimize the cross-entropy loss function for tokens $\bm{x}=(x_{1},\ldots,x_{T})$ drawn from this corpus:
$\min_{\bm{W}\in\mathbb{R}^{n}}\mathbb{E}_{\bm{x}\sim\mu}\left[\mathcal{L}_{\text{CE}}(\bm{x};\bm{W})\right]=-\mathbb{E}_{\bm{x}\sim\mu}\left[\sum_{t=1}^{T}\log P(x_{t}\mid\bm{x}_{1:t-1};\bm{W})\right]$ (1)
@@ II-B Continual Pretraining
$\min_{\bm{W}\in\mathbb{R}^{n}}-\mathbb{E}_{\bm{x}\sim\sigma}\left[\sum_{t=1}^{T}\log P(x_{t}\mid\bm{x}_{1:t-1};\bm{W},\bm{W}_{0})\right].$ (2)
@@
II-C Catastrophic Forgetting
Ideally, one would hope that such domain adaptation comes at no penalty. However, generally, a penalty is incurred. By shifting the focus from the general token distribution $\mu$ of the training dataset $\mathcal{D}$ to a new distribution $\sigma$, the model risks forgetting previously acquired knowledge. This risk depends on the size of the model, the size of the domain-specific dataset, and how different $\sigma$ is from $\mu$ [20]. Addressing this issue is crucial; otherwise, one may end up with a telecom-knowledge proficient LLM but lose reasoning capabilities, coding abilities, and general English language understanding in the process. Later in this paper, specifically in Section VII, we will detail how we addressed this issue in our continual pretraining framework. With this in mind, the next step in our framework involves curating a comprehensive telecommunications corpus, which we will refer to as Tele-Data.
III Tele-Data
@@
TABLE I: Tele-Data division across categories.
@@
TABLE II: Datasets filtering statistics.
Figure 2: Raw cross-entropy loss.
Figure 3: Cleaned cross-entropy loss.
Our collection of telecommunications material comprises four main sources: (1) scientific papers from arXiv, (2) 3GPP standards, (3) Wikipedia articles related to telecommunications, and (4) telecommunications-related websites extracted from Common Crawl dumps. This diverse range of sources ensures broad coverage of telecommunications knowledge and is essential for enabling the transfer of expertise from one aspect of telecommunications covered in one source to another, as will be shown later in Section VIII-A. The detailed composition of our dataset is reported in Table I, with LLama-3 tokenizer used to provide token counts as an example.
III-A Arxiv
@@
To assess the quality of our filtering process, we randomly sampled 500 arXiv papers and manually annotated them to determine their relevance to the telecommunications and networking domain. To ensure a balance between relevant and irrelevant examples, we drew an equal number of samples from both the computer science and electrical engineering categories given that the latter is more likely to contain telecommunications material, while the former is more likely to include unrelated papers. We then evaluated both the precision and recall of the filtering process. As shown in Table II, the process demonstrates high recall and moderate precision. High recall is particularly important, as it indicates that our filtering process captures virtually all relevant telecommunications material. The moderate precision can be attributed to the inclusion of papers from related domains. For instance, some papers investigating cloud computing topics were included. Since these additional domains are not orthogonal to telecommunications, they can provide transferable knowledge that may enhance the language model’s adaptation to the telecommunications domain. All in all, this demonstrates the effectiveness of our filtering process.
Cleaning. Following curation, we implement a comprehensive cleaning process for the papers, which includes: (1) removing comments, (2) flattening LaTeX sources, (3) substituting user-defined macros with standard LaTeX commands, (4) eliminating LaTeX native commands, and (5) standardizing citation formats and removing markup changes. We also remove figures and tables to focus on inline text and equations. Details on this procedure are provided in Appendix A. To assess the effectiveness of our cleaning process, we randomly selected 500 arXiv papers and derived a cross-entropy-based approach to evaluate cleanliness as detailed below. Let $\mathcal{X}=(\bm{x}^{1},\bm{x}^{2},\ldots,\bm{x}^{N})$ and $\mathcal{Y}=(\bm{y}^{1},\bm{y}^{2},\ldots,\bm{y}^{M})$ denote the non-overlapping chunks of $T$ tokens from the raw and cleaned datasets, respectively, where $N$ and $M$ are the total number of chunks. Next, let us define the cross-entropy loss for any token index $t\in\{1,\ldots,T\}$ in any chunk of either the raw or cleaned data using GPT-2 [4] as follows
@@
$\mathcal{L}^{\text{clean}}_{\text{CE}}(y_{t}^{i}|\bm{y}_{1:t-1}^{i})=-\log P(y_{t}^{i}|y_{1:t-1}^{i};\bm{W}_{\text{GPT-2}}),\quad i=1,\ldots,M.$ (3)
GPT-2 was chosen because it was primarily trained on web content, excluding arXiv scientific papers. Therefore, the cross-entropy here serves as a proxy for how closely the text resembles generic online content rather than the more complex arXiv papers, thus providing an indication of the dataset’s cleanliness. With this in mind, we set $T$ to 1024, in line with GPT-2’s training data [4], and define the total set of cross-entropy losses for the raw and cleaned datasets as
@@
$\mathcal{C}=\bigcup_{i=1}^{M}\{\mathcal{L}_{\text{CE}}(y_{t}^{i}|\bm{y}_{1:t-1}^{i}):t\in[1,1024]\},$ (4)